If(x, y, z, w) =0 then find (delx/dely) *(dely/delz) *(delz/delw) *(delw/delx)



1
Expert's answer
2021-01-18T15:33:28-0500

Let "F(x,y,z,w)=0, w=f(x,y,z)"

then:

"\frac{\partial x}{\partial y}=\frac{\partial y}{\partial z}=0"


So:

"\frac{\partial x}{\partial y}\frac{\partial y}{\partial z}\frac{\partial z}{\partial w}\frac{\partial w}{\partial x}=0"